# Fairness-aware Contextual Dynamic Pricing with Strategic Buyers

### v1

@@ 3.1 Buyers’ Fairness Perception and Learning Process
Given $t$ training samples of the form $(\bm{x}_{i},G_{i},p_{i})$, with $p_{i}$ the price set by the seller, the offline regression oracle OffRegP returns a prediction $\widehat{p}_{j}:\mathcal{X}\rightarrow\mathbb{R}^{+}$ for $j=0,1$. For any $10<\eta_{t}<$ 1, with probability at least $1-\eta_{t}$, we have
@@
where the offline learning guarantee $\mathcal{E}_{\mathcal{P},\eta_{t}}(t)$ is a function that decreases to 00 as $t\rightarrow\infty$.
The offline learning guarantee $\mathcal{E}_{\mathcal{P},\eta_{t}}(t)$ bounds the absolute distance between $\widehat{p}_{j}(\bm{x})$ and $p_{j}(\bm{x})$. Assumption 3 facilitates the flexibility of buyers’ learning methods and serves as a link between the buyer’s estimation error and the regret analysis. In general, $\mathcal{P}$ can be be any parametric or nonparametric function class. For instance, when $\mathcal{P}$ is the linear function class, the offline regression oracles can achieve the estimation error rate $O(\sqrt{t^{-1}\log(1/\eta_{t}}))$ (Lattimore and Szepesvári, 2020).
@@ 3.2 Fair Pricing Policy
6:     The seller sets a price $p_{t}\sim$ Unif$(0,B)$.
7:     The seller receives a demand $y_{t}$.
8:     The seller releases $(p_{t},G_{t},\bm{x}_{t})$ the public.
9:  end for
10:  Denote $\widetilde{\bm{x}}_{t}=(1,\bm{x}_{t}^{\top})^{\top}$ and the seller updates the parameter estimate, for $j\in\{0,1\}$,
$\widehat{\alpha}_{j},\widehat{\bm{\beta}}_{j}=\mathop{\arg\min}_{\alpha,\bm{\beta}}\sum_{t=1}^{T_{0}}(y_{t}-\alpha p_{t}-\bm{\beta}^{\top}\widetilde{\bm{x}}_{t})^{2}.$ (7)
11:  Exploitation Phase (Fairness-aware Optimal Pricing Policy):
12:  for $t=T_{0}+1,...,T$ do
13:     The buyer with $\bm{x}_{t}$ and $G_{t}\in\{0,1\}$ comes to the platform.
14:     The buyer learns the price difference and reveals the group status $G_{t}^{\prime}$ according to (6).
15:     The seller offers the price $p_{t}=p_{G^{\prime}}(\bm{x}_{t})$ with
@@
16:     The seller receives a demand $y_{t}$.
17:     The seller releases $(p_{t},G_{t}^{\prime},\bm{x}_{t})$ to the public.
18:  end for
@@
Exploration Phase. During this phase, the seller announces that a uniform pricing policy is implemented. At each time $t$, a buyer with $\bm{x}_{t}$ and $G_{t}$ enters the market. The seller provides the same price $p_{t}\sim$ Unif$(0,B)$ for both groups. In this case, buyers lack incentives to modify their private types and hence will reveal the true group status $G_{t}$ (Harris et al., 2022; Liu et al., 2024). After observing $p_{t}$, the buyer decides the demand $y_{t}$. The seller collects data $(\bm{x}_{t},G_{t},y_{t},p_{t})$. At the end of the exploration phase, the seller use the sale dataset $\{(\bm{x}_{t},G_{t},y_{t},p_{t})\}_{t=1}^{T_{0}}$, to estimate $\bm{\theta}_{j}=(\alpha_{j},\bm{\beta}_{j}^{\top})^{\top}$ for $j\in\{0,1\}$, see (7).
@@ 4 Theoretical Analysis
Suppose that Assumptions 1, 2 and 3 hold. At time $t$ in the exploitation phase, buyers learn the price difference $\widehat{\delta}$ using the offline regression oracle OffRegP. There exists a positive constant $c$ such that, for $t>c$, with probability at least $1-2\eta_{t-1}$, we have $\widehat{\delta}\leq C_{0},$ where $C_{0}$ represents manipulation cost of the group status.
Now, we provide some intuitive explanations on why our algorithm is able to prevent buyers’ strategic behaviors. Lemma 2 assures that $\widehat{\delta}\leq C_{0}$ holds with high probability. By (6), buyers are deterred from manipulation, as the cost of manipulation outweighs the benefits from manipulation.
@@
Let the assumptions of Lemma 2 hold. There exist positive constants $c_{1},c_{2},c_{3},c_{4}$ and $c_{5}$ such that when $T>c_{1}$, the total expected regret of the proposed pricing policy over the time horizon $T$ satisfies
$Regret_{T}<\sqrt{\left[\frac{c_{2}(d+2)+c_{3}}{q(1-q)}\right]T}+(c_{4}+c_{5}q)H(T),$
where $H(T)=\sum_{t=1}^{T}\eta_{t-1}$ with $\eta_{t}$ defined in Assumption 3.
@@
Under the assumptions of Theorem 2, if $\eta_{t}\leq\frac{1}{\sqrt{t}}$ when $t$ exceeding a certain constant, the total expected regret of the proposed pricing policy over the time horizon $T$ satisfies $Regret_{T}=O(\sqrt{T})$.
Corollary 1 follows directly from Theorem 2 under the condition $\eta_{t}\leq\frac{1}{\sqrt{t}}$ using the fact $\sum_{t=1}^{T}\frac{1}{\sqrt{t}}=O(\sqrt{T})$. This condition is attainable when buyers employ specific learning methods to learn the prices. For example, if buyers learn the price using the neural network, the condition $\eta_{t}\leq\frac{1}{\sqrt{t}}$ is met, with the offline learning guarantee $\mathcal{E}_{\mathcal{P},\eta_{t}}(t)$ decreasing over time (Ban et al., 2022). In Sections 5, we explore different learning methods used by buyers to assess our policy.
@@
Next, we provide an intuitive explanation for proving Theorem 3. The detailed proof is in Section S.6 of the appendix. The lower bound is established by constructing an uninformative price (Broder and Rusmevichientong, 2012; Xu and Wang, 2021; Fan et al., 2024). A price is deemed uninformative because all demand curves intersect at a common price, and no policy can gain information about the demand parameter. We choose the true demand parameter $\alpha$ as $\alpha_{0}=-2/5$. Using (3), we derive the optimal prices for group 0 and group 1 under the fairness constraint, denoted as $p_{0}^{*}(\alpha_{0})=1$ and $p_{1}^{*}(\alpha_{0})=3/4$, respectively. We then find that $\mathbb{E}(y_{t}|0,p_{0}^{*}(\alpha_{0}))=1/2$ and $\mathbb{E}(y_{t}|1,p_{1}^{*}(\alpha_{0}))=1/2$, indicating that all demand curves intersect at the optimal prices when the underlying parameter is $\alpha_{0}$. These optimal prices provide no information on the estimation of the demand parameter. We demonstrate that if a policy tries to learn model parameters, it must set prices away from the uninformative prices $p_{0}^{*}(\alpha_{0})$ and $p_{1}^{*}(\alpha_{0})$, thereby incurring large regret when the underlying parameter is $\alpha_{0}$. Furthermore, we establish that any policy failing to accurately learn the demand parameter $\alpha$ must also incur a cost in regret. Combining these facts, we can prove that any fair pricing policy achieves a regret lower bound $\Omega(\sqrt{T})$ in the setting presented in Theorem 3.
@@ 4.1 Outline of the Proof of Theorem 2
$p_{1t}=\left\{\begin{aligned} &-\frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}},&\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}\leq\delta-c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}},\\ &\widehat{\bm{\gamma}}_{1}^{\top}\widetilde{\bm{x}}_{t}-\delta+\widehat{\gamma}_{2},&\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}\geq\delta+c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}}.\end{aligned}\right.$
@@
$p_{0t}^{\prime}=\left\{\begin{aligned} &-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}},\ &\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}\leq\delta-c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}},\\ &\widehat{\bm{\gamma}}_{1}^{\top}\widetilde{\bm{x}}_{t}+\widehat{\gamma}_{2},\ &\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}\geq\delta+c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}}.\end{aligned}\right.$
The regret of our policy depends on the probability $\mathbb{P}(\widehat{\delta}\leq C_{0})$. We define the historical information up to time $t$ as $\mathcal{H}_{t}=\{\bm{x}_{1},\cdots,\bm{x}_{t},y_{1},\cdots,y_{t},p_{1},\cdots,p_{t}\}$. We also define ~$\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}\cup\{\bm{x}_{t+1}\}$ as the filtration including the feature $\bm{x}_{t+1}$. Given that the true group status at time $t$ is unknown, the expected regret at time $t$ in the exploitation phase is
@@
We can prove that $J_{1}$ is upper bounded by $\|\widehat{\bm{\theta}}_{0}-\bm{\theta}_{0}\|_{2}^{2}+\|\widehat{\bm{\theta}}_{1}-\bm{\theta}_{1}\|_{2}^{2}$. Noting that the number of samples from the exploration phase to estimate $\bm{\theta}_{0}$ and $\bm{\theta}_{0}$ is $T_{0}$, we can prove $J_{1}=O(1/T_{0})$ using Lemma 1. To establish an upper bound on $J_{2}$, we leverage Lemma 2 to upper bound $\mathbb{P}(\widehat{\delta}>C_{0})$. We can prove $J_{2}=O(\eta_{t-1})$. The expected regret at time $t$ in the exploitation phase is
@@
$Regret_{T}=O(T_{0})+O\left(\frac{T-T_{0}}{T_{0}}\right)+O\left(\sum_{t=1}^{T}\eta_{t-1}\right)=O(\sqrt{T})+O\left(\sum_{t=1}^{T}\eta_{t-1}\right),$
@@ 5 Simulation Study
In this section, we implement simulation studies to demonstrate the effectiveness of our policy. In the experiments (except for Section 5.2), we set the feature dimension as $d=3$. The demand parameters for buyers from group 0 is $\alpha_{0}=-1,\bm{\beta}_{0}=(2,1/2,1,1)^{\top}$. The demand parameters for buyers from group 1 is $\alpha_{1}=-1,\bm{\beta}_{1}=(1,1/4,1/2,1/2)^{\top}$. Assume the cost of manipulation is $C_{0}=0.8$. The seller would naturally select $\delta<C_{0}$ to discourage strategic behavior. In the experiments, the fairness level in the pricing policy is set to $\delta=0.799$. We denote the feature vector as $\bm{x}_{t}=(x_{1t},x_{2t},x_{3t})^{\top}$ and the features $x_{1t},x_{2t}$ and $x_{3t}$ are all i.i.d. from Unif(-2, 2). The noise distribution is $\epsilon_{t}\sim N(0,1)$.
As previous online pricing policies with fairness constraint (Xu et al., 2023; Chen et al., 2023a, b; Cohen et al., 2024) do not take strategic behaviors into consideration, they are not applicable in our problem. For the comparison, we consider the pricing policy without buyers’ fairness learning process mentioned in Section 2.1 as a benchmark fair pricing policy. In this policy, the seller provides prices with fairness constraints to strategic buyers while the buyers do not learn the fairness level. Without the fairness perception, the buyers always act strategically.

### v2

@@ 3.1 Buyers’ Fairness Perception and Learning Process
Given $t$ training samples of the form $(\bm{x}_{i},G_{i}^{\prime},p_{i})$, with $p_{i}$ the price set by the seller, the offline regression oracle OffRegP returns a prediction $\widehat{p}_{j}:\mathcal{X}\rightarrow\mathbb{R}^{+}$ for $j=0,1$. For some $\eta_{t}\in(0,1)$, with probability at least $1-\eta_{t}$, for all $\bm{x}$ from $\mathcal{X}$, we have
@@
where the offline learning guarantee $\mathcal{E}_{\mathcal{P},\eta_{t}}(t)$ is a function that decreases to 0 as $t\rightarrow\infty$.
The offline learning guarantee $\mathcal{E}_{\mathcal{P},\eta_{t}}(t)$ bounds the absolute distance between $\widehat{p}_{j}(\bm{x})$ and $p_{j}(\bm{x})$. Assumption 3 facilitates the flexibility of buyers’ learning methods and serves as a link between the buyer’s estimation error and the regret analysis. The predictor $\widehat{p}_{j}$ is assumed to come from a general function class $\mathcal{P}$, which may include neural networks, decision trees, or other supervised learning models. Because we do not assume which specific algorithm is used by buyers, the exact form of the estimation error $\mathcal{E}_{\mathcal{P},\eta_{t}}$ is left unspecified and may vary depending on the choice of learning algorithm. Theoretically, the realizability ($p_{i}(\cdot)\in\mathcal{P}$) is required. This condition ensures that the true pricing functions lie within the hypothesis class used by the buyer for learning. Under this assumption, the error $\mathcal{E}_{\mathcal{P},\eta_{t}}(t)$ in Assumption 3 can be guaranteed to decrease to 0 as $t\rightarrow\infty$. Without realizability, the error would include a non-vanishing approximation error term, and $\mathcal{E}_{\mathcal{P},\eta_{t}}(t)$ may not decrease to 0, which would undermine the regret analysis. Assumption 3 requires that the estimation error of the learned price should be small for all $\bm{x}$ from a fixed distribution. This assumption is reasonable because the covariates are sampled i.i.d. from a fixed distribution during both exploration and exploitation, ensuring that the coverage of the feature space expands as more data are collected. The only potential source of covariate shift comes from buyers’ reported group status in the exploitation phase. However, this concern diminishes over time, since the whole training dataset combines observations from both phases. In particular, data from the exploration phase are drawn i.i.d., which ensures diversity and helps mitigate covariate shift.
@@ 3.2 Fair Pricing Policy
6:  The buyer reveals $G_{t}^{\prime}=G_{t}$.
7:  The seller sets a price $p_{t}\sim$ Unif(0,B)(0,B).
8:  The seller receives a demand $y_{t}$.
9:  The seller releases $(p_{t},G_{t},\bm{x}_{t})$ the public.
10:end for
11: Denote $\widetilde{\bm{x}}_{t}=(1,\bm{x}_{t}^{\top})^{\top}$ and the seller updates the parameter estimate, for $j\in\{0,1\}$,
$\widehat{\alpha}_{j},\widehat{\bm{\beta}}_{j}=\mathop{\arg\min}_{\alpha\in\mathcal{A},\bm{\beta}\in\mathcal{B}}\sum_{t=1}^{T_{0}}\mathbb{I}(G_{t}=j)(y_{t}-\alpha p_{t}-\bm{\beta}^{\top}\widetilde{\bm{x}}_{t})^{2}.$ (7)
12:Exploitation Phase (Fairness-aware Optimal Pricing Policy):
13:for $t=T_{0}+1,...,T$ do
14:  The buyer with $\bm{x}_{t}$ and group status $G_{t}\in\{0,1\}$ comes to the platform.
15:  The buyer learns the price difference and reveals the group status $G_{t}^{\prime}$ according to (6).
16:  The seller offers the price $p_{t}=p_{G^{\prime}}(\bm{x}_{t})$ with
@@
17:  The seller receives a demand $y_{t}$.
18:  The seller releases $(p_{t},G_{t}^{\prime},\bm{x}_{t})$ to the public.
19:end for
@@
Exploration Phase. During this phase, the seller announces that a uniform pricing policy is implemented. At each time $t$, a buyer with $\bm{x}_{t}$ and $G_{t}$ enters the market. The seller provides the same price $p_{t}\sim$ Unif(0,B)(0,B) for both groups. In this case, buyers lack incentives to modify their private types and hence will reveal the true group status $G_{t}$ (Harris et al., 2022b; Liu et al., 2025). After observing $p_{t}$, the buyer decides the demand $y_{t}$. The seller collects data $(\bm{x}_{t},G_{t},y_{t},p_{t})$. At the end of the exploration phase, the seller use the sale dataset $\{(\bm{x}_{t},G_{t},y_{t},p_{t})\}_{t=1}^{T_{0}}$, to estimate $\bm{\theta}_{j}=(\alpha_{j},\bm{\beta}_{j}^{\top})^{\top}$ for $j\in\{0,1\}$, see (7).
@@ 4 Theoretical Analysis
Suppose that Assumptions 1, 2 and 3 hold. We assume $\delta<C_{0}$ where $C_{0}>0$ represents manipulation cost of the group status. At time $t$ in the exploitation phase, buyers learn the price difference $\widehat{\delta}$ using the offline regression oracle OffRegP. There exists a positive constant $c$ such that, for $t>c$, with probability at least $1-2\eta_{t-1}$, we have $\widehat{\delta}\leq C_{0}$.
We provide some intuitions on why our policy is able to prevent buyers’ strategic behaviors. Lemma 2 assures that $\widehat{\delta}\leq C_{0}$ holds with high probability. By (6), buyers are deterred from manipulation, as the manipulation cost outweighs the benefits from manipulation.
@@
Let the assumptions of Lemma 2 hold. There exist positive constants $c_{1},c_{2}$ and $c_{3}$ such that when $T>c_{1}$, the total expected regret of the proposed pricing policy over the time horizon $T$ satisfies
$Regret_{T}\leq\sqrt{\frac{c_{2}dT}{q(1-q)}}+c_{3}qH(T),$
where $H(T)=\sum_{t=1}^{T}\eta_{t}$ with $\eta_{t}$ defined in Assumption 3.
@@
Under the assumptions of Theorem 2, if $\eta_{t}\leq\frac{1}{\sqrt{t}}$ when $t$ exceeds a certain constant, the total expected regret of the proposed pricing policy over the time horizon $T$ satisfies $Regret_{T}=O(\sqrt{T})$.
Corollary 1 follows from Theorem 2 under the condition $\eta_{t}\leq\frac{1}{\sqrt{t}}$ using the fact $\sum_{t=1}^{T}\frac{1}{\sqrt{t}}=O(\sqrt{T})$. This condition is attainable when buyers employ specific methods to learn the prices. For example, if buyers learn the price using the neural network, the condition $\eta_{t}\leq\frac{1}{\sqrt{t}}$ is met, with the offline learning guarantee $\mathcal{E}_{\mathcal{P},\eta_{t}}(t)$ decreasing over time (Ban et al., 2022). In Sections 5, we explore different learning methods used by buyers to assess our policy.
@@
Next, we provide an intuitive explanation for proving Theorem 3. The detailed proof is in Section S.9.6 of the appendix. The lower bound is established by constructing an uninformative price (Broder and Rusmevichientong, 2012; Xu and Wang, 2021; Fan et al., 2024). A price is deemed uninformative because all demand curves intersect at a common price, and no policy can gain information about the demand parameter. We choose the true demand parameter $\alpha$ as $\alpha_{0}=-2/5$. Using (3), we derive the optimal prices for group 0 and group 1 under the fairness constraint, denoted as $p_{0}^{*}(\alpha_{0})=1$ and $p_{1}^{*}(\alpha_{0})=3/4$, respectively. We then find that $\mathbb{E}(y_{t}|0,p_{0}^{*}(\alpha_{0}))=1/2$ and $\mathbb{E}(y_{t}|1,p_{1}^{*}(\alpha_{0}))=1/2$, indicating that all demand curves intersect at the optimal prices when the underlying parameter is $\alpha_{0}$. These optimal prices provide no information on the estimation of the demand parameter. We demonstrate that if a policy tries to learn model parameters, it must set prices away from the uninformative prices $p_{0}^{*}(\alpha_{0})$ and $p_{1}^{*}(\alpha_{0})$, thereby incurring large regret when the underlying parameter is $\alpha_{0}$. Furthermore, we establish that any policy failing to accurately learn the demand parameter $\alpha$ must also incur a cost in regret. Combining these facts, we can prove that any fair pricing policy achieves a regret lower bound $\Omega(\sqrt{T})$ in the setting presented in Theorem 3.
@@ 4.1 Outline of the Proof of Theorem 2
$p_{1t}=\left\{\begin{aligned} &-\frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}},&\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}\leq\delta-c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}},\\ &\widehat{\bm{\gamma}}_{1}^{\top}\widetilde{\bm{x}}_{t}-\delta+\widehat{\gamma}_{2},&\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}>\delta+c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}}.\end{aligned}\right$.
@@
$p_{0t}^{\prime}=\left\{\begin{aligned} &-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}},\ &\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}\leq\delta-c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}},\\ &\widehat{\bm{\gamma}}_{1}^{\top}\widetilde{\bm{x}}_{t}+\widehat{\gamma}_{2},\ &\text{if}\ \ \frac{\widehat{\bm{\beta}}_{1}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{1}}-\frac{\widehat{\bm{\beta}}_{0}^{\top}\widetilde{\bm{x}}_{t}}{2\widehat{\alpha}_{0}}>\delta+c_{\delta}\sqrt{\frac{\log T_{0}}{T_{0}}}.\end{aligned}\right$.
The regret of our policy depends on the probability $\mathbb{P}(\widehat{\delta}\leq C_{0})$. We define the historical information up to time $t$ as $\mathcal{H}_{t}=\{\bm{x}_{1},\cdots,\bm{x}_{t},G_{1}^{\prime},\cdots,G_{t+1}^{\prime},y_{1},\cdots,y_{t},p_{1},\cdots,p_{t}\}$. We also define $\widetilde{\mathcal{H}}_{t}=\mathcal{H}_{t}\cup\{\bm{x}_{t+1},G_{t}^{\prime}\}$ as the filtration including $\bm{x}_{t+1}$ and $G_{t+1}^{\prime}$. Given that the true group status at time $t$ is unknown, the expected regret at time $t$ in the exploitation phase is
@@
We can prove that $J_{1}$ is upper bounded by $\|\widehat{\bm{\theta}}_{0}-\bm{\theta}_{0}\|_{2}^{2}+\|\widehat{\bm{\theta}}_{1}-\bm{\theta}_{1}\|_{2}^{2}$. Noting that the number of samples from the exploration phase to estimate $\bm{\theta}_{0}$ and $\bm{\theta}_{1}$ is $T_{0}$, we can prove $J_{1}=O(1/T_{0})$ using Lemma 1. To establish an upper bound on $J_{2}$, we leverage Lemma 2 to upper bound $\mathbb{P}(\widehat{\delta}>C_{0})$. We can prove $J_{2}=O(\eta_{t-1})$. The expected regret at time $t$ in the exploitation phase is
@@
$Regret_{T}=O(T_{0})+O\left(\frac{T-T_{0}}{T_{0}}\right)+O\left(\sum_{t=1}^{T}\eta_{t}\right)=O(\sqrt{T})+O\left(\sum_{t=1}^{T}\eta_{t}\right),$
@@ 5 Simulation Study
In this section, we implement simulation studies to demonstrate the effectiveness of our policy. In the experiments (except for Section 5.2), we set the feature dimension as $d=3$. The demand parameters for buyers from group 0 is $\alpha_{0}=-1,\bm{\beta}_{0}=(2,1/2,1,1)^{\top}$. The demand parameters for buyers from group 1 is $\alpha_{1}=-1,\bm{\beta}_{1}=(1,1/4,1/2,1/2)^{\top}$. Assume the cost of manipulation is $C_{0}=0.8$. The seller would naturally select $\delta<C_{0}$ to discourage strategic behavior. In the experiments, the fairness level in the pricing policy is set to $\delta=0.799$. We denote the feature vector as $\bm{x}_{t}=(x_{1t},x_{2t},x_{3t})^{\top}$ and the features $x_{1t},x_{2t}$ and $x_{3t}$ are all i.i.d. from Unif(-2, 2). The noise distribution is $\epsilon_{t}\sim N(0,1)$. Due to space limitations, additional simulations, including sensitivity analyses on hyperparameter choices and misspecified demand models, are provided in the appendix.
As previous online pricing policies with fairness constraint (Xu et al., 2023; Chen et al., 2026b, a; Cohen et al., 2025) do not take strategic behaviors into consideration, they are not applicable in our problem. For the comparison, we consider the pricing policy without buyers’ fairness learning process mentioned in Section 2.1 as a benchmark policy. In this policy, the seller provides prices with fairness constraints to strategic buyers while the buyers do not learn the fairness level. Without the fairness perception, the buyers always act strategically.